5)Janet drove her car due west for 3 hours while April drove her car due east for 2 hours. Janet drove 20km/hr faster than April. Their total distance driven combined is 400 km.
a) Complete the following table
------(km/hr)Distance-----Speed---------Time(h)
Janet----------------------j------------------
April----------------------a------------------

b) set up a system of linear equations ready to enter into your Ti-83 to represent this situation.


c)Find each of their speeds.


Please can you do this question and show the work. I tend to understand more when i see the steps that were taken to get the answer. Thank you! : D

Answer by stanbon(48568)

Janet drove her car due west for 3 hours while April drove her car due east for 2 hours. Janet drove 20km/hr faster than April. Their total distance driven combined is 400 km.
a) Complete the following table
Janet DATA:
time = 3 hrs ; rate = j km/hr ; distance = 3j km
----------------
April DATA:
time = 2 hrs ; rate = a km/hr ; distance = 2a km
--------------------------------
b) set up a system of linear equations ready to enter into your Ti-83 to represent this situation.
Rate equation: j-a = 20
distance eq:::3j+2a= 400
------------

c)Find each of their speeds.
Multiply thru the rate equation by 3:
3j-3a = 60
Subtract that from the distance equation and solve for "a":
5a = 340
a = 68 km/hr (April's rate)
-----------
Since j-a = 20, j = 88 km/hr (Janet's rate)
------------------------------------------------------------------
Cheers,
Stan H.